can you slove 5x-2y=-20
-4x+y=7?
please.

Answer by checkley77(12844)

5x-2y=-20
-4x+y=7 or y=4x+7 replace y in the first equation with (4x+7) & solve for x.
5x-2(4x+7)=-20
5x-8x-14=-20
-3x=-20+14
-3x=-6
x=-6/-3
x=2 ans.
y=4*2+7
y=8=7
y=15 ans.
Proof:
5*2-2*15=-20
10-30=-20
-20=-20